Aqueous ammonia is considered a weak base since it deprotonated water to a small extent to produce hydroxide ions.

Calculate the pH of a 2.0 mol/dm3 solution of Aqueous ammonia (Kb for ammonia =1.8*10^-5)


1
Expert's answer
2021-07-28T01:41:01-0400

Ammonia is a weak base, so we use formula:

[OH^-] = √Kb × Molarity

= √1.8×10^-5 × 2 M

= √3.6 × 10^-5

= 6 × 10^-3

pOH = - log[OH^-]

= - log6 × 10^-3

= 3 - log6

pH = 14 - pOH

= 14 - (3-log6)

= 10 + log6

= 11.778